# Decoding Futures Price Dynamics: A Regularized Sparse Autoencoder for Interpretable Multi-Horizon Forecasting and Factor Discovery

### v3

@@ Abstract
Commodity futures price volatility creates significant economic challenges, necessitating accurate multi-horizon forecasting. Predicting these prices is complicated by diverse interacting factors (macroeconomic, supply/demand, geopolitical). Current models often lack transparency, limiting strategic use. This paper presents a Regularized Sparse Autoencoder (RSAE), a deep learning framework for simultaneous multi-horizon commodity futures prediction and discovery of interpretable latent market drivers. The RSAE forecasts prices at multiple horizons (e.g., 1-day, 1-week, 1-month) using multivariate time series. A key L1 regularization ($\|\mathbf{z}\|_{1}$) on its latent vector $\mathbf{z}$ enforces sparsity, promoting parsimonious explanations of market dynamics through learned factors representing underlying drivers (e.g., demand shifts, supply shocks). Drawing from energy-based models and sparse coding, the RSAE optimizes predictive accuracy while learning sparse representations. Evaluated on historical Copper and Crude Oil futures data with numerous indicators, our findings suggest the RSAE offers competitive multi-horizon forecasting accuracy and data-driven insights into price dynamics via its interpretable latent space, a notable advantage over traditional black-box approaches.
@@ 1 Introduction
Given these limitations, this research addresses the question: How can we develop a deep learning framework that achieves high predictive accuracy for commodity futures prices across multiple horizons and simultaneously extracts a sparse, interpretable set of latent factors representing key market drivers? We propose a Regularized Sparse Autoencoder (RSAE) framework that utilizes an autoencoder backbone [99] with a time-lagged predictive objective and L1 sparsity regularization [97] on its latent code. This design, inspired by energy-based models [61] and sparse coding [75], forces the model to explain market dynamics through a parsimonious set of learned factors.
Our contribution lies in the specific design and application of this RSAE framework to multi-horizon commodity futures forecasting, aiming to enhance both accuracy and transparency. The remainder of this paper details the related work, the RSAE methodology, data, experimental setup, results, and discussion.
@@ 2.4 Key Machine Learning Techniques Applied
Note: Strengths, weaknesses, and representative studies are derived from the provided research material.
@@
3 Methodology: The Regularized Sparse Autoencoder (RSAE) Framework
This section details the Regularized Sparse Autoencoder (RSAE) framework, an approach specifically designed for the dual task of multi-horizon commodity futures price prediction and the discovery of interpretable latent market factors. The methodology draws inspiration from energy-based models [61] and sparse coding [75], employing an explicit inference step to determine a sparse latent representation that best explains future price movements.
@@ 3.1 Problem Formulation
We consider a set of $d$ market features at time $t$, $\mathbf{x}_{t}\in\mathbb{R}^{d}$. The model input is a look-back window $X=X_{t-W+1:t}=(\mathbf{x}_{t-W+1},\ldots,\mathbf{x}_{t})\in\mathbb{R}^{W\times d}$. Our objective is to predict the futures price $p_{t}$ at $N$ horizons $\tau_{1}<\ldots<\tau_{N}$, with target $Y=Y_{t}=(p_{t+\tau_{1}},\ldots,p_{t+\tau_{N}})^{\top}\in\mathbb{R}^{N}$. The RSAE learns a mapping from $X$ to $Y$ via a sparse latent variable $\mathbf{z}\in\mathbb{R}^{m}$ ($m\ll W\times d$).
3.2 Model Architecture
The RSAE comprises three neural network modules (Figure 1).
Figure 1: Conceptual architecture of the Regularized Sparse Autoencoder (RSAE) framework.
@@ 3.2 Model Architecture
    History Predictor (Pred): $\bm{h}=\text{Pred}(X;\theta_{\text{Pred}})$, maps $X$ to a context vector $\bm{h}\in\mathbb{R}^{h_{\text{dim}}}$, using LSTMs [48], GRUs [22], or 1D CNNs [7].
@@
    Decoder (Dec): $\tilde{Y}=\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}})$, predicts future prices $\tilde{Y}\in\mathbb{R}^{N}$ from $\bm{z}$ and $\bm{h}$.
@@
    Encoder (Enc) (for training): $\hat{\bm{z}}=\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$, performs amortized inference of $\bm{z}^{*}$ from actual future $Y$ and $\bm{h}$ [34, 57]. A separate encoder, $\text{Enc}_{\text{XtoZ}}(X;\theta_{\text{XtoZ}})$, is trained to predict $\bm{z}^{*}$ from $X$ alone for test-time forecasting. $\text{Enc}_{\text{XtoZ}}$ is trained with a mean-squared error objective to regress the K-step optimal z*; its parameters are not updated through the prediction loss.
3.3 Latent Variable Framework: Energy Function and Inference
We define an energy function $E(Y,\bm{h},\bm{z})=C(Y,\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}}))+\lambda R(\bm{z})$ [61]. $C(Y,\tilde{Y})=\frac{1}{N}\|Y-\tilde{Y}\|_{2}^{2}$ is the MSE prediction cost [12]. $R(\bm{z})=\|\bm{z}\|_{1}$ is the L1-norm sparsity regularizer [97]. The optimal latent variable $\bm{z}^{*}=\arg\min_{\bm{z}}E(Y,\bm{h},\bm{z})$ is found via $K$-step (sub)gradient descent [15].
3.3.1 Selection and Convergence of Inference Steps (K)
The number of iterative steps, $K$, in the inference process for $\bm{z}^{*}$ is a hyperparameter balancing thoroughness of energy minimization against computational cost. To determine an appropriate value for $K$, preliminary experiments were conducted on the validation set. This involved tracking the average energy function value, $E(Y,\bm{h},\bm{z})$, over a range of potential iterations (e.g., $k=1$ to $k=15$) for multiple representative batches of validation samples.
The observed convergence behavior is depicted in Figure 2. The mean energy (averaged across validation batches) typically exhibited a sharp decrease during the initial iterations (approximately $k=1$ to $k=7$). Beyond this initial phase, the rate of energy reduction substantially diminished, with the curve beginning to plateau around $k=8$ to $k=10$. While formal convergence to a global minimum is not guaranteed for non-convex energy landscapes, the iterative process aims to find a suitable local minimum that provides a sparse and predictive latent code. Based on the empirical observation that iterations beyond $K=10$ provided only marginal decreases in the objective energy at an increased computational cost, $K=10$ was selected for all subsequent experiments in this study. A more extensive analysis of convergence properties under varying conditions remains an avenue for future work.
Figure 2: Mean objective energy function value $E(Y,\bm{h},\bm{z})$ versus the number of inference iterations ($k$), averaged over representative validation set batches. The plot demonstrates typical convergence behavior, with energy decreasing rapidly initially before plateauing. The selected $K=10$ (dashed red line) represents a point of diminishing returns for energy reduction.
3.4 Test-Time Inference
At test time, future $Y$ is unknown. We train an additional inference encoder, $\text{Enc}_{\text{XtoZ}}(X;\theta_{\text{XtoZ}})$, supervised by the $\bm{z}^{*}$ values obtained during the training phase (from the $K$-step optimization using known $Y_{\text{train}}$). Thus, for a new test input $X_{\text{test}}$, the latent code is obtained directly: $\bm{z}_{\text{test}}=\text{Enc}_{\text{XtoZ}}(X_{\text{test}};\theta_{\text{XtoZ}})$. This amortized inference from $X$ alone is used for generating forecasts.
3.5 Training Objective and Parameter Updates
The primary training objective minimizes the prediction loss using the $K$-step inferred $\bm{z}^{*}$: $L_{\text{Pred}}^{*}=C(Y,\text{Dec}(\bm{z}^{*},\bm{h};\theta_{\text{Dec}}))$. Parameters $\theta_{\text{Pred}}$ and $\theta_{\text{Dec}}$ are updated based on gradients from $L_{\text{Pred}}^{*}$. If the training-time Encoder $\text{Enc}(Y,\bm{h})$ is used, it is trained with an inference loss $L_{\text{Inf}}=\|\bm{z}^{*}-\text{Enc}(Y,\bm{h})\|_{2}^{2}$. The test-time encoder $\text{Enc}_{\text{XtoZ}}(X)$ is trained separately to predict $\bm{z}^{*}$ from $X$.
3.6 Overall Training Algorithm
@@
    Initialize $\theta_{\text{Pred}},\theta_{\text{Dec}},\theta_{\text{XtoZ}}$.
@@
    For each training batch $\{(X^{(j)},Y^{(j)})\}$:
    1. (a)
      $\bm{h}^{(j)}=\text{Pred}(X^{(j)})$.
    2. (b)
      Infer $\bm{z}^{*(j)}$ via $K$-step minimization of $E(Y^{(j)},\bm{h}^{(j)},\bm{z})$.
    3. (c)
      $L_{\text{Pred}}^{*(j)}=C(Y^{(j)},\text{Dec}(\bm{z}^{*(j)},\bm{h}^{(j)}))$.
    4. (d)
      Update $\theta_{\text{Pred}},\theta_{\text{Dec}}$ via gradients of $\text{mean}(L_{\text{Pred}}^{*})$ (e.g., Adam [56]).
    5. (e)
      Update $\theta_{\text{XtoZ}}$ to predict $\bm{z}^{*(j)}$ from $X^{(j)}$.
@@
    Validate using $L_{\text{Pred}}^{*}$ (with $\bm{z}_{\text{val}}=\text{Enc}_{\text{XtoZ}}(X_{\text{val}})$) for tuning and early stopping [82].

### v4

# Sparse Latent Factor Forecaster (SLFF) with Iterative Inference for Transparent Multi-Horizon Commodity Futures Prediction

@@ Abstract
Commodity futures are volatile. Forecasting across horizons with interpretable drivers remains challenging. We propose the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), a structured-prediction latent-variable model that combines sparse coding, unrolled optimization, and amortized inference. SLFF explicitly optimizes a sparse latent code to explain multi-horizon futures trajectories and trains an encoder whose outputs are validated against the optimization-based solution before deployment. The method is paired with an information-set-aware pipeline (vintage macro releases, lag-aware fills, leakage checks) and evaluated under rolling-origin folds against representative statistical and neural baselines. We provide quantitative criteria for factor labeling and directional diagnostics that account for no-change regimes. On Copper and WTI futures (2005–2023), SLFF achieves competitive RMSE/MAE, improves directional skill beyond persistence, and yields factors that are stable across seeds and linked to measurable fundamentals. Code, diagnostics, and information-set specifications are released for reproducibility.
TITLE PAGE
Manuscript Title:
Sparse Latent Factor Forecaster (SLFF) with Iterative Inference for Transparent Multi-Horizon Commodity Futures Prediction
Manuscript Type:
Research Article
Author:
Abhijit Gupta, PhD
Affiliation: Independent Researcher
ORCID: 0000-0002-6292-3789
Email: abhijit.gupta_in@outlook.com
Corresponding Author:
Abhijit Gupta, PhD (email above)
Running Title (suggested):
SLFF for Commodity Futures Forecasting
Keywords (suggested):
commodity forecasting; multi-horizon prediction; latent factors; interpretability; time series; deep learning
Word Count (approx., excluding references):
7204
Figures:
4
Tables:
16
Funding:
None
Competing Interests:
None
Ethics Approval:
Not applicable (no human or animal subjects).
Data Availability:
See the manuscript Data Availability and Licensing section. The study uses a mix of public and proprietary sources; proprietary datasets require the corresponding licenses.
Permissions to Reproduce Material:
All figures and tables are original. Underlying data include public and proprietary sources; permissions/licenses are as described in the Data Availability section.
@@ 1 Introduction
Given these limitations, this research addresses the question: How can we develop a latent-variable forecasting framework that remains faithful to the information set, attains robust multi-horizon accuracy, and exposes stable economic factors? We answer with the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), a structured-prediction model that uses an explicit latent-factor optimization layer together with a LISTA-style amortized encoder [34]. The latent code $\mathbf{z}$ is never treated as a reconstruction target; it is the solution of a constrained regression problem that explains future trajectories under an L1 sparsity prior.
This paper makes the following contributions:
  1. 1.
    Method: We introduce SLFF, a sparse latent-factor forecaster with iterative inference, and describe a latent inference objective that supports both test-time amortization and (optionally) unrolled, differentiable optimization.
  2. 2.
    Information set: We specify an information-set-aware data pipeline, including release calendars and vintage macroeconomic series (ALFRED/FRED-MD, IMF WEO vintages), and provide an “available-at-tt” specification to prevent mixed-frequency leakage.
  3. 3.
    Evaluation: We evaluate on six rolling-origin folds (train: 6y, val: 1y, test: 1y) with test years spanning 2013–2023 (every two years), reporting mean $\pm$ standard deviation across folds and random seeds.
  4. 4.
    Baselines: We compare against PatchTST, Temporal Fusion Transformer (TFT), N-HiTS, DeepAR, and tuned gradient boosting models under matched hyperparameter budgets and learning-rate schedules.
  5. 5.
    Ablations: We present ablations over sparsity, iterative inference, latent dimensionality, and training regimes to quantify their impact on deployed forecasting accuracy and latent interpretability.
  6. 6.
    Directional diagnostics: We report directional metrics that account for no-change regimes (no-change rate, class-conditional hit rates, MCC, calibration curves, confusion matrices).
  7. 7.
    Factor validation: We evaluate factor stability via Procrustes alignment across random seeds and validate factors against observable disruption/event indicators, including counterfactual perturbation tests.
  8. 8.
    Reproducibility: We provide code (JAX/Flax), experiment manifests, Dockerized data pipelines, compute footprints, and an executable feature notebook to support replication and extension.
The remainder of this paper revisits the related work, describes SLFF in detail, specifies the curated information set, explains the experimental protocol, and presents empirical/interpretability evidence before concluding with limitations and future work.
@@ 2.4 Key Machine Learning Techniques Applied
Note: The table provides a compact overview; details and examples are given in the cited references.
@@
3 Methodology: Sparse Latent Factor Forecaster with Iterative Inference
This section details the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), an approach specifically designed for the dual task of multi-horizon commodity futures price prediction and the discovery of interpretable latent market factors. The methodology draws inspiration from structured prediction energy networks [55], sparse coding [69], and differentiable optimization layers [3], employing an explicit inference step to determine a sparse latent representation that best explains future price movements.
@@ 3.1 Problem Formulation
Let $\mathbf{x}_{t}\in\mathbb{R}^{d}$ denote the feature vector available at the market close of day $t$. We form an input tensor $X_{t}=(\mathbf{x}_{t-W+1},\ldots,\mathbf{x}_{t})\in\mathbb{R}^{W\times d}$ with look-back $W$. The goal is to predict a vector of future log prices $Y_{t}=(p_{t+\tau_{1}},\ldots,p_{t+\tau_{N}})^{\top}$ with horizons $\tau_{1}<\cdots<\tau_{N}$. Instead of mapping $X_{t}$ directly to $Y_{t}$, SLFF introduces a low-dimensional latent code $\mathbf{z}_{t}\in\mathbb{R}^{m}$ with $m\ll Wd$ that is inferred to best explain $Y_{t}$ subject to a sparsity prior. Forecasts are produced via $\tilde{Y}_{t}=\text{Dec}(\mathbf{z}_{t},\mathbf{h}_{t})$ where $\mathbf{h}_{t}$ summarizes history. The latent code is therefore a structured prediction variable, not a reconstruction target.
3.2 Model Components
The architecture (Figure 1) contains four modules:
@@ 3.2 Model Components
    History summarizer $\text{Pred}(X_{t})$: a two-layer LSTM (or GRU / temporal convolution) that compresses $X_{t}$ into $\mathbf{h}_{t}\in\mathbb{R}^{h_{\text{dim}}}$.
@@
    Decoder $\text{Dec}(\mathbf{z},\mathbf{h})$: a lightweight MLP that maps the latent code and context to the $N$ forecasting horizons. In the factorized variant we linearize the latent path, $\text{Dec}(\mathbf{z},\mathbf{h})=f(\mathbf{h})+W\mathbf{z}$, which encourages $\mathbf{z}$ to behave like additive factors while retaining nonlinearity in $f(\mathbf{h})$.
@@
    Iterative inference layer: a LISTA-like unrolled optimizer that solves for an optimization-refined latent $\mathbf{z}_{t}^{\star}$ using gradient steps on an energy function (Section 3.3). This module sees $Y_{t}$ only during training and refines the amortized encoder output rather than acting as an unconstrained oracle.
  4. 4.
    Amortized encoder $\text{Enc}_{X\rightarrow Z}(X_{t})$: a network trained to approximate $\mathbf{z}_{t}^{\star}$ from $X_{t}$ alone. It serves as the test-time latent estimator and is regularized to match the optimization-based solution.
Figure 1: Conceptual architecture of the SLFF framework. Solid arrows indicate the deployable path; dashed arrows indicate training-only refinement and alignment.
3.3 Latent Energy and Iterative Inference
We define
$E(Y_{t},\mathbf{h}_{t},\mathbf{z})$ $=C\big(Y_{t},\text{Dec}(\mathbf{z},\mathbf{h}_{t})\big)+\lambda\|\mathbf{z}\|_{1}+\mu\|\mathbf{z}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2},$ (1)
$C(Y_{t},\tilde{Y}_{t})$ $=\frac{1}{N}\|Y_{t}-\tilde{Y}_{t}\|_{2}^{2},$ (2)
where $\lambda$ enforces sparsity and $\mu$ ties inference to the amortized encoder. We denote the optimization-refined latent by
$\mathbf{z}^{\star}_{t}=\arg\min_{\mathbf{z}}E(Y_{t},\mathbf{h}_{t},\mathbf{z}),$ (3)
which is a constrained refinement of $\text{Enc}_{X\rightarrow Z}(X_{t})$. For clarity, the unconstrained oracle (not used in training or evaluation) corresponds to the special case $\mu=0$:
$\mathbf{z}^{\dagger}_{t}=\arg\min_{\mathbf{z}}\;C\big(Y_{t},\text{Dec}(\mathbf{z},\mathbf{h}_{t})\big)+\lambda\|\mathbf{z}\|_{1}.$ (4)
We solve the refinement with $K$ steps of proximal gradient descent,
$\mathbf{z}^{(k+1)}=\text{SoftThresh}_{\alpha\lambda}\big(\mathbf{z}^{(k)}-\alpha(\nabla_{\mathbf{z}}C(Y_{t},\text{Dec}(\mathbf{z}^{(k)},\mathbf{h}_{t}))+2\mu(\mathbf{z}^{(k)}-\text{Enc}_{X\rightarrow Z}(X_{t})))\big),$ (5)
where the shrinkage operator is applied element-wise. In practice we warm-start inference at $\mathbf{z}^{(0)}=\text{Enc}_{X\rightarrow Z}(X_{t})$ to encourage predictable latent codes; unless otherwise stated, this includes the $\mu=0$ setting, so the $\mu=0$ sweep row reflects a warm-started refinement rather than a cold-start oracle. In the $\mu$-sweep we tune $\alpha$ per $\mu$ to ensure stable convergence (monotone energy decrease), so the sweep is not confounded by step-size instability. The iterations are differentiable, enabling end-to-end training or truncated backpropagation. When enabled (not in the main experiments), the contrastive regularizer described in Section 5 is applied only during encoder training and does not modify the inference energy.
3.3.1 Choosing $K$
We sweep $K\in\{1,\ldots,15\}$ on validation folds, monitoring $E$ and reconstruction errors. Figure 2 shows that energy decreases rapidly until $k\approx 8$ and flattens thereafter. We therefore use $K=10$ for the default setting and expose ablations in Section 5.4. The trend is consistent across commodities and seeds.
Figure 2: Energy convergence of the iterative inference module averaged across validation folds. Error bars denote $\pm$ one standard deviation across random seeds.
3.4 Bridging Training and Test Time
During training, we compute $\mathbf{z}^{\star}_{t}$ using the iterative solver conditioned on the realized $Y_{t}$. At deployment, only $X_{t}$ is accessible; hence $\mathbf{z}_{t}=\text{Enc}_{X\rightarrow Z}(X_{t})$. To ensure that $\text{Enc}_{X\rightarrow Z}$ is not learning an impossible target, we:
  • minimize $\|\mathbf{z}^{\star}_{t}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$ in alternating updates with the forecasting loss (two-stage loop, with a stop-gradient on $\mathbf{z}^{\star}_{t}$);
  • optionally (not in main experiments) add a contrastive regularizer that shuffles $Y_{t}$ within the batch while holding $X_{t}$ fixed, forcing the encoder-matched latents to be invariant to inconsistent futures; this prevents $\text{Enc}_{X\rightarrow Z}$ from chasing sample-specific $Y$ noise that $\mu$ alone would still allow via the refinement path;
  • include the proximity term $\mu\|\mathbf{z}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$ in the inference objective and warm-start $\mathbf{z}^{(0)}=\text{Enc}_{X\rightarrow Z}(X_{t})$ to keep $\mathbf{z}^{\star}_{t}$ predictable from $X_{t}$;
  • treat $\text{Enc}_{X\rightarrow Z}(X_{t})$ as a stop-gradient inside the inference loop (default two-stage), so $\mathbf{z}^{\star}_{t}$ is computed with a frozen encoder snapshot and cached before encoder updates; this avoids a moving-target objective unless the unrolled end-to-end variant is explicitly enabled.
  • report alignment diagnostics (Table 10) to show that $R^{2}$ between $\mathbf{z}^{\star}_{t}$ and $\text{Enc}_{X\rightarrow Z}(X_{t})$ is around 0.73 on validation/test folds and that cosine similarity remains high on held-out splits.
Section 5.4 shows that removing this consistency training noticeably erodes both forecasting accuracy and interpretability.
3.5 Differentiable Variants
To address the concern that $\text{Enc}_{X\rightarrow Z}$ parameters were previously frozen, we now support three regimes:
@@ 3.5 Differentiable Variants
    Two-stage training (default, stop-grad inference): $\text{Enc}_{X\rightarrow Z}$ is treated as a stop-gradient inside the inference loop, so $\mathbf{z}^{\star}_{t}$ is computed with a fixed encoder snapshot. The encoder is then updated only via the match/contrast losses, not the forecast loss, which controls train-test mismatch and discourages a decoder that only works with optimization-refined latents.
@@
    Unrolled end-to-end (fully differentiable): We backpropagate through all $K$ inference steps and allow gradients to flow through $\text{Enc}_{X\rightarrow Z}$, akin to learned ISTA. This removes the stop-grad assumption and directly couples inference and prediction, but is $\sim 18\%$ slower; we include it in ablations for completeness.
@@
    Deployment fine-tune: after encoder alignment, we optionally fine-tune Pred and Dec using $\hat{\mathbf{z}}=\text{Enc}_{X\rightarrow Z}(X_{t})$ in the prediction loss to optimize the deployed path.
3.6 Training Objective and Algorithm
The total loss is
$L=L_{\text{pred}}+\beta L_{\text{match}}+\gamma L_{\text{contrast}},$ (6)
where $L_{\text{pred}}=C(Y_{t},\text{Dec}(\mathbf{z}^{\star}_{t},\mathbf{h}_{t}))$, $L_{\text{match}}=\|\mathbf{z}^{\star}_{t}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$, and the contrastive term is defined by shuffling $Y_{t}$ within a batch while holding $X_{t}$ fixed:
$Y_{t}^{\text{shuff}}$ $=\Pi(Y_{t}),\quad\mathbf{z}^{\star,\text{shuff}}_{t}=\arg\min_{\mathbf{z}}E(Y_{t}^{\text{shuff}},\mathbf{h}_{t},\mathbf{z}),$ (7)
$L_{\text{contrast}}$ $=\mathbb{E}\left[\cos^{2}\!\left(\text{Enc}_{X\rightarrow Z}(X_{t}),\mathbf{z}^{\star,\text{shuff}}_{t}\right)\right],$ (8)
which penalizes alignment between encoder latents and inconsistent future refinements. This complements the $\mu$ proximity constraint by discouraging a latent code that overfits to accidental $Y$ noise while remaining close to the encoder output. Unless noted, we set $\gamma=0$ in the main results, so the contrastive term is a diagnostic option and is not used in the main experiments. Hyperparameter $\beta$ is fixed at 5 across folds/commodities; $\gamma$ is set to 0 unless the contrastive regularizer is explicitly enabled. Unless Stage 3 is enabled, Pred/Dec are optimized on $\mathbf{z}^{\star}$ while evaluation uses $\text{Enc}_{X\rightarrow Z}(X)$; Table 13 in Appendix A reports the resulting diagnostic gap. We keep $L_{\text{pred}}$ on $\mathbf{z}^{\star}$ to preserve the refinement learning signal; mixed-loss or deployment-path fine-tune variants are optional and not used in the main tables. In the optional deployment fine-tune we replace $L_{\text{pred}}$ with $C(Y_{t},\text{Dec}(\hat{\mathbf{z}}_{t},\mathbf{h}_{t}))$ where $\hat{\mathbf{z}}_{t}=\text{Enc}_{X\rightarrow Z}(X_{t})$. Unless stated otherwise, reported results use the two-stage model without deployment fine-tuning (Stage 3 omitted).
The training loop proceeds as follows.
  1. 1.
    Sample a batch $\{(X^{(j)},Y^{(j)})\}$.
  2. 2.
    Compute $\mathbf{h}^{(j)}=\text{Pred}(X^{(j)})$.
  3. 3.
    Run $K$ inference steps (warm-started at $\text{Enc}_{X\rightarrow Z}(X^{(j)})$) to obtain $\mathbf{z}^{\star(j)}$, treating $\text{Enc}_{X\rightarrow Z}$ as stop-gradient unless the unrolled end-to-end variant is used.